OpenStudy (anonymous):

How do I eliminate the parameter to find a Cartesian equation of the curve? (See problem below, please)

OpenStudy (anonymous):

\(\ \Huge x=t^2, y=6-3t \).

OpenStudy (anonymous):

here t=(6-y)/3 so x=((6-y)/5)^2

OpenStudy (anonymous):

So basically do you solve for t and then substitute that in to the next equation?

OpenStudy (amistre64):

solve t in terms of x, and sub it into the y part

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

@amistre64 So, in this case t=sqrt(x), so I get \(\ -3\sqrt{x}+6 ?\)

OpenStudy (anonymous):

y= that is.

OpenStudy (amistre64):

keep in mind the t^2 has 2 solutions; and that the true nature of the curve, if convention is followed, would result in 2 equations for y but yes

OpenStudy (anonymous):

2 Equations?

OpenStudy (amistre64):

\[t^2=x\to\ t=\pm\sqrt{x}\] sub into y= to get:\[y=6\pm3\sqrt{x}\]

OpenStudy (anonymous):

That makes sense. Thank you, @amistre64!

OpenStudy (amistre64):

youre welcome; matricks is fine to depending on how conventional they want the cartesian equation